# Comparing the information content of probabilistic representation spaces

### v2

@@ Appendix A Appendix: Extended channel similarity results
Appendix B Appendix: Information estimation using Bhattacharyya distinguishability matrices
@@
In this work, we assume that the data distribution can be approximated by the empirical distribution, $p(x)\approx\sum_{i}^{N}\delta(x-x_{i})/N$, simplifying Eqn. 5 so that $c_{i}\equiv 1/N$ and $H(U|C)=H(U|X)$ because the identity of the component is equivalent to the identity of the datum. Finally, we assume that the set of posterior distributions for a representative sample of size $M$ of the dataset, taken for the fingerprint, adequately approximates the empirical distribution. Larger samples may be necessary in different scenarios when the amount of information transmitted by channels is larger than a handful of bits, but $M=1000$ appeared sufficient for the analyses of this work.
The corrections to NMI and VI required the information conveyed by two measurements from different channels, $I(X;U,V)$ as well as from the same channel, $I(X;U,U^{\prime})$. The matrix of Bhattacharyya coefficients given measurements $U$ and $V$ is simply the elementwise product of the coefficients given $U$ and the coefficients given $V$. The posterior in the joint space of $U$ and $V$ is factorizable given $x$—i.e., $p(u,v|x)=p(u|x)p(v|x)$—because the stochasticity in each channel is independent. The same is true for the joint space of $U$ and $U^{\prime}$, two draws from the same channel. The Bhattacharyya coefficient of the joint variable simplifies,
@@
Appendix C Appendix: Broken triangle inequality for the generalized VI
@@
Appendix D Appendix: Are NMI and VI interchangeable?
@@
In Fig. 13, we compare NMI and VI (estimated via Bhattacharyya matrices) as similarity or distance measures for the pairwise comparisons between channels used in Fig. 3. Specifically, as measures of similarity, we plot NMI against $\exp(-\text{VI})$, and for distance we plot $-\log(\text{NMI})$ against VI. We find that NMI and VI are non-trivially related, shown clearly by the horizontal and vertical swaths of points where one of the two measures is roughly constant while the other varies considerably.
Interestingly, the corresponding NMI and VI comparisons in Fig. 13 show multiple distinct arcs of channel similarity, as well as clear vertical bands where NMI has discerning power and VI does not.
Appendix E Appendix: Implementation specifics
We have uploaded code to the following repository: https://github.com/murphyka/representation-space-info-comparison.
@@ 4. Conv2DTranspose: 32 4$\times$4 ReLU kernels, stride 2, padding ‘same’
Monte Carlo mutual information estimation: We sampled random data points and then a random embedding vector from each posterior distribution. Then we computed the log ratio of the likelihood under the corresponding posterior and the aggregated posterior (iterating over the entire dataset), and repeated $N$ times to estimate the following expectation:
For the analysis of Fig. 4, $N=2\times 10^{5}$ for all datasets except dsprites and celebA, where $N=6\times 10^{4}$; the standard error of the estimate was on the order of 0.01 bits or less. The displayed points on the plots are weighted means of the values, weighted by the propagated uncertainty on each quantity; the error bars are the standard error of the weighted mean.

### v3

@@ Appendix A Appendix: Extended channel similarity results
Appendix B Appendix: Extended results on synthesized latent space comparison (Sec. 4.1)
We include in Fig. 11 the full set of comparisons between the Bhattacharyya fingerprint and the average predictions of a trained classification head for the nine synthetic latent spaces of Sec. 4.1. Fig. 12 shows the pairwise scatter plots for the methods of comparing the latent spaces.
Figure 11: Full comparison of Bhattacharyya fingerprints and classification outputs from latent spaces of Fig. 2.
Figure 12: Pairwise scatter plot comparison of similarity measures from Fig. 2.
Appendix C Appendix: Information estimation using Bhattacharyya distinguishability matrices and Monte Carlo estimator
@@
In this work, we assume that the data distribution can be approximated by the empirical distribution, $p(x)\approx\sum_{i}^{N}\delta(x-x_{i})/N$, simplifying Eqn. 6 so that $c_{i}\equiv 1/N$ and $H(U|C)=H(U|X)$ because the identity of the component is equivalent to the identity of the datum. Finally, we assume that the set of posterior distributions for a representative sample of size $M$ of the dataset, taken for the fingerprint, adequately approximates the empirical distribution. Larger samples may be necessary in different scenarios when the amount of information transmitted by channels is larger than a handful of bits, but $M=1000$ appeared sufficient for the analyses of this work.
The generalizations of NMI and VI required the information conveyed by two measurements from different channels, $I(X;U,V)$ as well as from the same channel, $I(X;U,U^{\prime})$. The matrix of Bhattacharyya coefficients given measurements $U$ and $V$ is simply the elementwise product of the coefficients given $U$ and the coefficients given $V$. The posterior in the joint space of $U$ and $V$ is factorizable given $x$—i.e., $p(u,v|x)=p(u|x)p(v|x)$—because the stochasticity in each channel is independent. The same is true for the joint space of $U$ and $U^{\prime}$, two draws from the same channel. The Bhattacharyya coefficient of the joint variable simplifies,
@@
For the Monte Carlo estimator, we sampled random data points from the dataset $x\sim p(x)$ and then a random embedding vector from each posterior distribution, $u\sim p(u|x)$. Then we computed the log ratio of the likelihood under the corresponding posterior $p(u|x)$ and the aggregated posterior $p(u)$, which was gotten by iterating over the entire dataset. We repeated $N$ times to estimate the following expectation:
For the analysis of Fig. 4, $N=2\times 10^{5}$ for all datasets except dsprites and celebA, where $N=6\times 10^{4}$; the standard error of the estimate was on the order of 0.01 bits or less.
Figure 14: Partially aggregated posterior for the Monte Carlo estimator. For three models selected from those of Fig. 4, we seleced random subsets of the full dataset to use when computing the aggregated posterior for estimating the transmitted information $I(U;X)$ of the full 10-dimensional latent space. The points are offset horizontally by random amounts for visibility, and the error bars show the standard error on the mean of the Monte Carlo samples.
Computing the fully aggregated posterior is time consuming; could a subsample of the dataset be used instead? In Fig. 14, we varied the fraction of the dataset used to compute the (partially) aggregated posterior for a full 10-dimensional latent space for three models that transmit an intermediate amount of information about the dataset on which they were trained. Each point represents the mutual information estimate following Eqn. 8 for a random subset of the dataset, and the extent of its error bar is given by the standard error: the standard deviation of the set of values divided by the square root of the size of the set. We find that the transmitted information is relatively robust to using half of the dataset when computing the aggregated posterior, but deleterious effects grow when reducing the dataset by a factor of 10 or more. Interestingly, the dsprites estimate was largely insensitive to using 3% of the dataset, perhaps due to its highly structured nature.
In Fig. 4, the NMI and VI values are averaged over all pairwise comparisons for the five representation spaces, with weights given by the inverse squared uncertainty on each pairwise value. The uncertainty was calculated by propagating the uncertainty of Monte Carlo estimates of $I(X;U)$ according to the expression for VI or NMI. To be specific, as $\text{VI}(U,V)=2I(X;U,V)-I(X;U,U^{\prime})-I(X;V,V^{\prime})$, the propagated uncertainty is given by
$\Delta\text{VI}(U,V)=\sqrt{4\Delta I(X;U,V)^{2}-\Delta I(X;U,U^{\prime})^{2}-\Delta I(X;V,V^{\prime})^{2}},$ (9)
with $\Delta I(X;U,V)$ the standard error from the Monte Carlo estimate. The expression for the uncertainty on the NMI estimate contains many more terms, and we will not reproduce them here. Finally, the error bars in Fig. 4 are the standard error of the weighted mean,
$\Delta Q=\left(\sum_{i}^{n}(\Delta Q_{i})^{-2}\right)^{-\frac{1}{2}},$ (10)
for quantity $Q$.
Appendix D Appendix: Broken triangle inequality for the generalized VI
@@
Appendix E Appendix: Are NMI and VI interchangeable?
@@
In Fig. 16, we compare NMI and VI (estimated via Bhattacharyya matrices) as similarity or distance measures for the pairwise comparisons between channels used in Fig. 3. Specifically, as measures of similarity, we plot NMI against $\exp(-\text{VI})$, and for distance we plot $-\log(\text{NMI})$ against VI. We find that NMI and VI are non-trivially related, shown clearly by the horizontal and vertical swaths of points where one of the two measures is roughly constant while the other varies considerably. Interestingly, the corresponding NMI and VI comparisons in Fig. 16 show multiple distinct arcs of channel similarity, as well as clear vertical bands where NMI has discerning power and VI does not.
Appendix F Appendix: Implementation specifics
Code to reproduce the experiments of Sec. 4 can be found at the following repository: https://github.com/murphyka/representation-space-info-comparison. The heart of the codebase is in utils.py, containing the Bhattacharyya and Monte Carlo calculations of $I(U;X)$ and the NMI/VI calculations.